# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

### v1

@@ B.2 Quantum Upper Bound of Set-Inclusion Problem
We restate and prove Lemma 3.10 as follows: See 3.10
@@
The proof of Lemma 3.10 is similar to the proof of Lemma 3.9. The only essential difference is that 3.11 is replaced by its quantum speedup version, A.18. For simplicity, we discuss the following two cases as the proof of Lemma 3.9. For other cases, we can obtain the same result similarly. In the following proof, for $x\in\{0,1\}^{n}$, let $S_{x}\coloneqq\left\{i:x_{i}=1\right\}$. Let $\Pi\coloneqq\sum_{i\in S_{x\land y}}\ket{i}\bra{i}$ |.
@@
    $n_{1}=c,n_{2}=a-c$. In this case, the problem to compute $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ can be reduced equivalently to estimate $\frac{|x\land y|}{|x|}$. We first prove some special unitary operations can be constructed with $O(\log n)$ qubits communication.
@@
      Since Alice knows all information about $x$, she can implement $S_{x}$ such that $\mathcal{S}_{x}\ket{0}=\frac{1}{\sqrt{a}}\sum_{i\in S_{x}}\ket{i}$ by herself without communication. Let |$\ket{\phi_{1}}\coloneqq\mathcal{S}_{x}\ket{0}$. Then we have
      $\bra{\phi_{1}}\Pi\ket{\phi_{1}}=\frac{|x\land y|}{|x|}.$
@@
      $O_{x}\ket{i}$ $\coloneqq(-1)^{x_{i}}\ket{i},$
      $O_{y}\ket{i}$ $\coloneqq(-1)^{y_{i}}\ket{i},$
      $O_{x\land y}\ket{i}$ $\coloneqq(-1)^{x_{i}y_{i}}\ket{i}.$
      Then $I-2\Pi=O_{x\land y}$. For any |$\ket{\phi}\coloneqq\sum_{i\in[n]}\alpha_{i}\ket{i}$, Alice and Bob can perform $I-2\Pi$ to any state |$\ket{\phi}$ using $O(\log n)$ qubits of quantum communication as follows: i) Alice performs $O_{x}$ to |$\ket{\phi}$ and sends |$\ket{\phi^{\prime}}=O_{x}\ket{\phi}=\sum_{i\in[n]}\alpha_{i}(-1)^{x_{i}}\ket{i}$ to Bob; ii) Bob performs $O_{y}$ to |$\ket{\phi^{\prime}}$ and obtains $\sum_{i\in[n]}\alpha_{i}(-1)^{x_{i}y_{i}}\ket{i}=O_{x\land y}\ket{\phi}$.
    By A.18, a quantum speedup version of 3.11, and using an argument similar to Lemma 3.9, we can obtain the following conclusion: using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$ unitary operations $\mathcal{S}_{x}$ and $I-2\Pi$, Alice and Bob can estimate $\frac{|x\land y|}{|x|}$ with errors at most $O\left(\frac{g}{a}\right)$. Furthermore, they can decide whether $|x\land y|\geq c+g$ or $|x\land y|\leq c-g$ with success probability at least $1-1/(6\log n)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits communication.
@@
    $n_{1}=c$, $n_{2}=n-a-b+c$ and $a+b<n$. In this case, the problem to compute $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ can be reduced equivalently to estimate $\frac{|x\land y|}{|\overline{x}\oplus y|}$. Similar to Case 1, we first prove Alice and Bob can construct a quantum sampler $\mathcal{S}_{\overline{x}\oplus y}$ such that
    $\mathcal{S}_{\overline{x}\oplus y}\ket{0}\ket{0}=\frac{1}{\sqrt{|N|}}\sum_{f\in[N]}\ket{ff}\ket{f(x,y)}=\ket{\phi_{2}},$
    using $O\left(\log n\right)$ qubits communication, where $N=C\cdot\frac{n_{1}n_{2}}{g^{2}}$ for some enough large constant $C$ and each $f$ encodes a random function such that $f(x,y)$ is a uniform distribution of the elements in $S_{\overline{x}\oplus y}$. Specifically, the construction procedure is as follows: First, Alice and Bob use $O\left(\log\frac{n_{1}n_{2}}{g^{2}}\right)$ qubits communication to transform |$\ket{0}$ into a maximum entanglement bipartite state $\frac{1}{\sqrt{N}}\sum_{f\in[N]}\ket{ff}$. Moreover, both Alice and Bob hold one part of the state. Second, by Fact 3.12, Alice and Bob can sample an element from $S_{\overline{x}\oplus y}$ uniformly using $O(\log n)$ bits of communication. Equivalently, Alice and Bob can generate a random function $f$ using public coins, and output $f(x,y)$ using $O(\log n)$ bits of communication such that the distribution of the output is a uniform distribution of the elements in $S_{\overline{x}\oplus y}$. Since quantum circuits can simulate classical circuits efficiently, Alice and Bob can perform a uniform operation that transforms |$\ket{ff}\ket{b}$ into |$\ket{ff}\ket{f(x,y)+b}$ for any $f\in N$ using $O\left(\log n\right)$ qubits communication, where both Alice and Bob hold one part of |$\ket{ff}$, $b\in\left\{0,1,...,n-1\right\}$ and the addition is with modulo $n$. As a whole, Alice and Bob can construct the above quantum sampler $\mathcal{S}_{\overline{x}\oplus y}$ using $O\left(\log\frac{n_{1}n_{2}}{g^{2}}+\log n\right)=O\left(\log n\right)$ qubits communication. Furthermore, since $f(x,y)$ is a uniform distribution of the elements in $S_{\overline{x}\oplus y}$ for any $f\in\left\{0,1,...,N-1\right\}$, $N=C\cdot\frac{n_{1}n_{2}}{g^{2}}$ for some enough large constant $C$, and
    $\bra{\phi_{2}}\left(I\otimes\Pi\right)\ket{\phi_{2}}=\frac{|\left\{f:f(x,y)\in S_{x\land y}\right\}|}{N},$
@@
    $|\bra{\phi_{2}}\left(I\otimes\Pi\right)\ket{\phi_{2}}-\frac{|x\land y|}{|\overline{x}\oplus y|}|=O\left(\frac{g}{m}\right)$ (9)
@@
By A.18, a quantum speedup version of 3.11, and using an argument similar to Lemma 3.9, we can obtain the following conclusion, using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$ unitary operations $\mathcal{S}_{|\overline{x}\oplus y|}$ and $I-2\Pi$, Alice and Bob can estimate $\bra{\phi_{2}}\left(I\otimes\Pi\right)\ket{\phi_{2}}$ with errors at most $O\left(\frac{g}{m}\right)$. By Equation (9), Alice and Bob also can estimate $\frac{|x\land y|}{|\overline{x}\oplus y|}$ with errors at most $O\left(\frac{g}{m}\right)$. Next, using the same argument as Lemma 3.9, we can prove that Alice and Bob can decide whether $|x\land y|\geq c+g$ or $|x\land y|\leq c-g$ with success probability at least $1-1/(6\log n)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits communication. ∎
Appendix C Log-Rank Conjecture of Permutation-Invariant Functions
In this section, we restate and prove Lemma 4.2. See 4.2
Proof.
$\log\binom{n}{a}\leq\min\left\{a,n-a\right\}\cdot\log n.$ (12)
Combining Equations 10, 11 and 12, the communication cost of the protocol is
Appendix D Log-Approximate-Rank Conjecture of Permutation-Invariant Functions
@@
    Then
    $\mathsf{ESetInc}_{a,b,c,g}^{n}\left(x,y\right)=\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\left(x^{\prime},y^{\prime}\right),$
    Thus, $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is a submatrix of $\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$.
@@
    $\mathsf{ESetInc}_{a,b,c,g}^{n}(x,y)=\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}(x,\overline{y})=\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}(\overline{x},y).$
@@
    Since $\mathsf{ESetInc}_{a,b,c,g}^{n}(x,y)=\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}(\underbrace{x\cdots x}_{k},\underbrace{y\cdots y}_{k})$, we have $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is a submatrix of $\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}$.
@@
Let $P$ be the $\left(2k,k,f_{k,l}\right)$-pattern matrix, where $f_{k,l}$ is defined in Equation 7. By Lemma 3.8, $P$ is a submatrix of $\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}$. Thus, we have
$\log\left(\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)\geq\log\left(\operatorname{\widetilde{rank}}\left(P\right)\right).$
By Facts A.7 and A.15 and Equation 8, we have
@@
$\log\left(\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)=\Omega\left(\operatorname{\mathsf{\widetilde{deg}}}\left(f_{k,l}\right)\right)=\Omega\left(\sqrt{kl}\right),$
where the second equality comes from Equation 8. ∎

### v2

@@ B.2 Quantum Upper Bound of Set-Inclusion Problem
We restate and prove Lemma 3.9 as follows: See 3.9
@@
The proof of Lemma 3.9 is similar to the proof of Lemma 3.8. The only essential difference is that 3.10 is replaced by its quantum speedup version, A.18. For simplicity, we discuss the following two cases as the proof of Lemma 3.8. For other cases, we can obtain the same result similarly. In the following proof, for $x\in\{0,1\}^{n}$, let $S_{x}\coloneqq\left\{i:x_{i}=1\right\}$. Let $|\Pi\coloneqq\sum_{i\in S_{x\land y}}\ket{i}\bra{i}$.
@@
    $n_{1}=c,n_{2}=a-c$. In this case, the problem to compute $\mathrm{SetInc}_{a,b,c,g}^{n}(x,y)$ can be reduced equivalently to estimate $\frac{|x\land y|}{|x|}$. We first prove some special unitary operations can be constructed with $O(\log n)$ qubits communication.
@@
[…]
      $|.\bra{\phi_{1}}\Pi\ket{\phi_{1}}=\frac{|x\land y|}{|x|}$.
@@
[…]
    By A.18, a quantum speedup version of 3.10, and using an argument similar to Lemma 3.8, we can obtain the following conclusion: using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$ unitary operations $\mathcal{S}_{x}$ and $I-2\Pi$, Alice and Bob can estimate $\frac{|x\land y|}{|x|}$ with errors at most $O\left(\frac{g}{a}\right)$. Furthermore, they can decide whether $|x\land y|\geq c+g$ or $|x\land y|\leq c-g$ with success probability at least $1-1/(6\log n)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits communication.
@@
    $n_{1}=c$, $n_{2}=n-a-b+c$ and $a+b<n$. In this case, the problem to compute $\mathrm{SetInc}_{a,b,c,g}^{n}(x,y)$ can be reduced equivalently to estimate $\frac{|x\land y|}{|\overline{x}\oplus y|}$. Similar to Case 1, we first prove Alice and Bob can construct a quantum sampler $\mathcal{S}_{\overline{x}\oplus y}$ such that
[…]
@@
[…]
@@
[…]
Appendix C Log-Approximate-Rank Conjecture of Permutation-Invariant Functions
@@
    Then $\mathrm{ESetInc}_{a,b,c,g}^{n}\left(x,y\right)=\mathrm{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\left(x^{\prime},y^{\prime}\right)$. Thus, $\mathrm{ESetInc}_{a,b,c,g}^{n}$ is a submatrix of $\mathrm{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$.
@@
    $\mathrm{ESetInc}_{a,b,c,g}^{n}(x,y)$ $=\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}(x,\overline{y})$
    $=\mathrm{ESetInc}_{n-a,b,b-c,g}^{n}(\overline{x},y).$
@@
    Since $\mathrm{ESetInc}_{a,b,c,g}^{n}(x,y)=\mathrm{ESetInc}_{ka,kb,kc,kg}^{kn}(\underbrace{x\cdots x}_{k},\underbrace{y\cdots y}_{k})$, we have $\mathrm{ESetInc}_{a,b,c,g}^{n}$ is a submatrix of $\mathrm{ESetInc}_{ka,kb,kc,kg}^{kn}$.
@@
Let $P$ be the $\left(2k,k,f_{k,l}\right)$-pattern matrix, where $f_{k,l}$ is defined in (10). By Lemma 3.7, $P$ is a submatrix of $\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}$. Thus, we have
$\log\left(\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)\geq\log\left(\operatorname{\widetilde{rank}}\left(P\right)\right).$
By Facts A.7 and A.15 and (11), we have
@@
$\log\left(\operatorname{\widetilde{rank}}\left(\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)$ $=\Omega\left(\operatorname{\mathsf{\widetilde{deg}}}\left(f_{k,l}\right)\right)$
$=\Omega\left(\sqrt{kl}\right),$
where the second equality comes from (11). ∎
Appendix D Communciation Complexity of Gap-Hamming-Distance Problem
Ref. [23] gave the lower and upper bounds on the randomized communication complexity of Gap-Hamming-Distance as Lemmas D.1 and D.3. We show Lemmas D.1 and D.3 can be expressed as Lemmas D.2 and D.4 equivalently.
Lemma D.1 (Lemma 3.3 in [23]).
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming distances of $\Delta(x,y)$ when $|x|=a,|y|=b$. Then
$R(\mathrm{GHD}_{a,b,c,g}^{n})$ $=\Omega\left(\frac{\min\left\{a,b,c,n-a,n-b,n-c\right\}}{g}\right),$ (13)
$R(\mathrm{GHD}_{a,b,c,g}^{n})$ $=\Omega\left(\log\left(\frac{\min\left\{c,n-c\right\}}{g}\right)\right).$
Lemma D.2.
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming weights of $|x\land y|$ when $|x|=a,|y|=b$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. Then $R(\mathsf{SetInc}_{a,b,c,g}^{n})$ has two lower bounds: $\Omega\left(\frac{n_{2}}{g}\right)$ and
$\Omega\left(\log\frac{\min\left\{a+b-2c,n-a-b+2c\right\}}{g}\right).$
Proof.
By Definitions 1.12 and 1.13, $\mathsf{SetInc}_{a,b,c,d}^{n}$ and $\mathrm{GHD}_{a,b,a+b-2c,2g}^{n}$ are the same problems. Thus, we have
$R(\mathsf{SetInc}_{a,b,c,d}^{n})=R(\mathrm{GHD}_{a,b,a+b-2c,2g}^{n}).$
By Lemma D.1, $R(\mathrm{GHD}_{a,b,a+b-2c,2g}^{n})$ has two lower bounds:
$\Omega\left(\frac{\min\left\{a,b,a+b-2c,n-a,n-b,n-(a+b-2c)\right\}}{g}\right),\\$
and
$\Omega\left(\log\frac{\min\left\{a+b-2c,n-a-b+2c\right\}}{g}\right).$
Since
$a$ $=(a-c)+c,$
$b$ $=(b-c)+c,$
$a+b-2c$ $=(a-c)+(b-c),$
$n-a$ $=(n-a-b+c)+(b-c),$
$n-b$ $=(n-a-b+c)+(a-c),$
$n-(a+b-2c)$ $=(n-a-b+c)+c,$
and $n_{1},n_{2}$ are smallest two numbers in ac,c,bc,nab+ca-c,c,b-c,n-a-b+c, we have
$\min\left\{a,b,a+b-2c,n-a,n-b,n-(a+b-2c)\right\}$
$=n_{1}+n_{2}$
$\geq n_{2}.$
Lemma D.3 (Lemma 3.4 in [23]).
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming distances of $\Delta(x,y)$ when $|x|=a,|y|=b$. If $a\leq b\leq n/2$, then $R(\mathrm{GHD}_{a,b,c,g}^{n})$ is
$O\left(\min\left\{\left(\frac{a}{g}\right)^{2}\log\left(\frac{b}{g}\right),\left(\frac{c}{g}\right)^{2},\left(\frac{n-c}{g}\right)^{2}\right\}\right).$ (14)
Lemma D.4.
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming weights of $|x\land y|$ when $|x|=a,|y|=b$. If $a\leq b\leq n/2$, then
$R(\mathsf{SetInc}_{a,b,c,g}^{n})=O\left(\left(\frac{n_{2}}{g}\right)^{2}\log\frac{b}{g}\right).$
Proof.
Since $a\leq b\leq n/2$, similar to the proof of Lemma D.2, we have
$\min\left\{a,a+b-2c,n-(a+b-2c)\right\}$
$=\min\left\{a,b,a+b-2c,n-a,n-b,n-(a+b-2c)\right\}$
$=n_{1}+n_{2}$
$\leq 2n_{2}.$
By Lemma D.3, we have
$R(\mathsf{SetInc}_{a,b,c,d}^{n})$
$=R(\mathrm{GHD}_{a,b,a+b-2c,2g}^{n})$
$=O\left(\frac{\min\left\{a,a+b-2c,n-(a+b-2c)\right\}^{2}}{(2g)^{2}}\log\left(\frac{b}{2g}\right)\right)$
$=O\left(\left(\frac{n_{2}}{g}\right)^{2}\log\left(\frac{b}{g}\right)\right).$
Acknowledgement
The authors would like to thank the editor and anonymous reviewers for their helpful comments and suggestions. We also would like to express our gratitute to Arkadey Chattopadhyay for drawing our attention to the proper attribution of the resolution of the Log-Approximate-Rank Conjecture for $\mathrm{XOR}$-symmetric functions. Accordingly, we have revised the manuscript to correctly acknowledge the work of Chattopadhyay and Mande [18], who first established this result.